OpenStudy (anonymous):

I need someone to walk me through how to get the answer to this: (5x-3y^5)^2

OpenStudy (anonymous):

"What is a simpler form of the product?"

OpenStudy (gorv):

open the square (a+b)^2=a^2+b^2+2ab

OpenStudy (gorv):

(x^n)^2=x^2n

OpenStudy (gorv):

@babysmith baby can u do it now

OpenStudy (anonymous):

No I don't understand that.

OpenStudy (gorv):

can u open the square????

OpenStudy (gorv):

@babysmith

OpenStudy (anonymous):

I don't know what 'open the square' means?

OpenStudy (skullpatrol):

Have you heard of FOIL?

OpenStudy (anonymous):

First, outside, inside, last.

OpenStudy (skullpatrol):

Yep, use it :)

OpenStudy (anonymous):

But the problem doesn't look like a problem I would use FOIL on.

OpenStudy (skullpatrol):

$$ (5x-3y^5)^2=(5x-3y^5)(5x-3y^5)$$

OpenStudy (skullpatrol):

Or we could use the distributive property if you want?

OpenStudy (anonymous):

I have the answer, by the way, I'm just not sure how to get to the answer.

OpenStudy (skullpatrol):

$$a(b-c)=ab-ac$$ $$(5x-3y^5)(5x-3y^5)=(5x-3y^5)(5x)-(5x-3y^5) (3y^5)$$ $$(b-c)a=ba-ca$$ $$(5x)(5x)-(3y^5)(5x)-[(5x)(3y^5)-(3y^5)(3y^5)]$$ $$=25x^2-15xy^5-15xy^5+9y^{10}$$ $$=25x^2-30xy^5+9y^{10}$$

OpenStudy (anonymous):

For some reason I still don't understand. Thank you for trying anyway.

OpenStudy (skullpatrol):